A clinical trial tests a method designed to increase the probability of conceiving a girl. In the study 604 babies were​ born, and 302 of them were girls. Use the sample data to construct a 99​% confidence interval estimate of the percentage of girls born. Based on the​ result, does the method appear to be​ effective?
Answer #1

Solution :

Given that,

n = 604

x = 302

Point estimate = sample proportion = \hat p = x / n = 302/604=0.5

1 - \hat p   = 1-0.5 =0.5

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96   ( Using z table )

Margin of error = E = Z\alpha / 2    * \sqrt (((\hat p * (1 - \hat p )) / n)

= 1.96 (\sqrt((0.5*0.5) /604 )

= 0.0399

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.5-0.0399 < p < 0.5+0.0399

0.4601< p < 0.5399

The 95% confidence interval for the population proportion p is : 0.4601, 0.5399